A window consists of a rectangle surmounted by semi-circle. What shape gives the most area for a given perimeter?


1
Expert's answer
2021-01-19T02:21:09-0500

Let us skerch a picture of such a window:




Let "a" and "b" be lengths of a corresponding rectangle. Then the surmounted semi-circle is of radius "\frac{a}{2}".

The perimeter of the figure is "P=2b+a+\frac{\pi a}{2}=2b+\frac{2+\pi}{2}a" and the area is "S=ab+\frac{1}{2}\pi(\frac{a}{2})^2=ab+\frac{\pi a^2}{8}." It follows that "b=\frac{P}{2}-\frac{2+\pi}{4}a" and therefore, "S=a(\frac{P}{2}-\frac{2+\pi}{4}a)+\frac{\pi a^2}{8}=\frac{P}{2}a-\frac{2+\pi}{4}a^2+\frac{\pi a^2}{8}
=\frac{P}{2}a-\frac{4+\pi}{8}a^2". Since the perimeter is constant, "S'=\frac{P}{2}-\frac{4+\pi}{4}a". If "S'=0," then "\frac{P}{2}-\frac{4+\pi}{4}a=0", and thus "a=\frac{2P}{4+\pi}". Taking into account that "S''=-\frac{4+\pi}{4}<0", we conclude that "a=\frac{2P}{4+\pi}" is a point of maximum of "S". It follows that "b=\frac{P}{2}-\frac{2+\pi}{4}\frac{2P}{4+\pi}=\frac{P}{2}-\frac{(2+\pi)P}{2(4+\pi)}=\frac{4P+\pi P-2P-\pi P}{2(4+\pi)}=\frac{P}{4+\pi}", and thus "a=2b". We conclude that for "b=\frac{P}{4+\pi}" and "a=2b=\frac{2P}{4+\pi}" the area is maximum.